OpenStudy (anonymous):

help help help for a medal the company discovered that it costs $16 to produce 2 widgets, $18 to produce 4 widgets, and $48to produce 10 widgets. Using the quadratic function, find the cost of producing 6 widgets.

OpenStudy (anonymous):

there is actually a way to solve this but it is rather a pain i would use technology http://www.wolframalpha.com/input/?i=quadratic+%282%2C16%29%2C%284%2C18%29%2C%2810%2C48%29

OpenStudy (anonymous):

looks like your quadratic function is \[f(x)=\frac{1}{2}x^2-2x+18\]

OpenStudy (anonymous):

to answer the question posed, find \[f(6)\]

OpenStudy (anonymous):

This problem looks awfully familiar.

OpenStudy (anonymous):

@satellite73 so would i just plug in 6 into the x's?

OpenStudy (anonymous):

have you done one similar? @Johnbc

OpenStudy (anonymous):

yes that is what you would do

OpenStudy (anonymous):

I have watched a similar problem be performed but in a different manner

OpenStudy (anonymous):

you have to solve a three by three system of equations to do this, and it is annoying

OpenStudy (anonymous):

oh okay well thanks for the help website helped out great @satellite73

OpenStudy (anonymous):

yw note the syntax

OpenStudy (anonymous):

seems really difficult ... @Johnbc

OpenStudy (anonymous):

Yeah that is exactly right, it turns out to be 3 different values we are looking for (a, b, and c) and there will be 3 different equations so it will be system of equations but I like and trust satellite73's method.

OpenStudy (anonymous):

Quite simple but tedious.

OpenStudy (anonymous):

would you be able to explain or is it to late ? @Johnbc

OpenStudy (anonymous):

\[ax^2+bx+c\] \[a\times 2^2+b\times 2+c=16\] or \[4a+2b+c=16\] is equation one

OpenStudy (anonymous):

similarly get two more equations solve the system

OpenStudy (anonymous):

Of course, As a quadratic function \[Cost = ax^2+bx+c\] Where X is the amount of widgets. So for each \[16 = a(2)^2+b(2)+c\] \[18 = a(4)^2+b(4)+c\] \[48 = a(10)^2+b(10)+c\]

OpenStudy (anonymous):

You can use one the equations to solve for the values of a, b, and c as system.

OpenStudy (anonymous):

In the method satellite showed, you would plug in 6

OpenStudy (anonymous):

thank you so much i got it my answer was 24 thank you :) @Johnbc

OpenStudy (anonymous):

My pleasure

OpenStudy (anonymous):

by any chance would you be able to figure out the quadratic function ? like satellite73

OpenStudy (anonymous):

I am going to solve it using the other method so you can refer to it next time if you get stuck again. Using \[16= 4a+2b+c\] We subtract it from the other 2 equations so \[18=16a+4b+c\]\[-(16 = 4a+2b+c)\]_________________________ 2 = 12a + 2b And\[48 = 100a+10b+c\]\[-(16=4a+2b+c)\]______________________________ 32 = 96a + 8b

OpenStudy (anonymous):

Using both our new equations, I multiply my smaller equation by 8 to get 16 = 96a + 16b Because we want to eliminate another term as we did for the "c" terms so this eliminates the "a", \[32 = 96a + 8b\]\[-(16= 96a +16b)\]____________________ 16 = 0 - 8b b = -2 Using b =-2 back into 2 = 12a +8b \[2 = 12a +2(-2) \]\[2 = 12a-4\]\[6 = 12a\]\[a = \frac{ 6 }{ 12 } = \frac{ 1 }{ 2 } \]

OpenStudy (anonymous):

We know have b = -2 and a = 1/2 so using our original equation \[16= 4a +2b +c\] we can solve for c \[16 = 4(\frac{ 1 }{ 2 })+2(-2) +c\] \[16 = 2 -4 +c\] \[18 = c\]

OpenStudy (anonymous):

Now that we have a. b, and c we can find our quadratic function as \[cost = \frac{ 1 }{ 2 }x^2-2x+18\]

OpenStudy (anonymous):

ommmmgg thank you so muchh !

OpenStudy (anonymous):

My pleasure

OpenStudy (anonymous):

It is exactly as Satellite73 said it would be! & Now you have the steps for reference.